Can someone please help me!! It would be greatly appreciated:
Find the center, vertices and foci of the hyperbola: y^2=x^2+8x

Answer by stanbon(48581)

Find the center, vertices and foci of the hyperbola: y^2=x^2+8x
---------------
Rewrite:
x^2+8x -y^2 = 0
Complete the square on the x-terms:
x^2+8x+16 - y^2 = 16
Factor:
(x+4)^2/16 - (y-0)^2/16 = 1
----------------------------
Center is at (-4,0)
a = 4 ; b = 4 ; c = sqrt(16 + 16) = 4sqrt2
-------------
Vertices:
(-4+4,0) or (0,0)
(-4-4,0) or (-8,0)
---------------
Foci:
(-4+4sqrt2,0)
(-4-4sqrt2,0)
=================
Cheers,
Stan H.